OpenStudy (anonymous):

Which expression is equivalent to sin2(π2−x)sinx?

zepdrix (zepdrix):

\[\Large\rm \sin^2(\pi^2-x)\sin x\]This is the expression?

OpenStudy (anonymous):

sin2(π2−x)sinx first part is right (x/2-x) divided by sin x

zepdrix (zepdrix):

\[\Large\rm \frac{\sin^2\left(\frac{\pi}{2}-x\right)}{\sin x}\]

zepdrix (zepdrix):

Sine and cosine are co-functions, meaning:\[\Large\rm \sin(90-x)=\cos x\]So our numerator simplifies,\[\Large\rm \frac{\sin^2\left(\frac{\pi}{2}-x\right)}{\sin x}=\frac{\cos^2x}{\sin x}\]

zepdrix (zepdrix):

Which can be simplified a lil bit more I guess.